OpenStudy (anonymous):

Given the following function, find: (a) vertex, (b) axis of symmetry, (c) intercepts, (d) domain, (e) range, (f) intervals where the function is increasing,(g) intervals where the function is decreasing, and (h) the graph of the function f(x)=-x^2-6

OpenStudy (campbell_st):

find the axis of symmetry x = -b/2a in this question, a = -1, b = 0 and c= -6 so the axis of symmetry is x= 0 or the y-axis vertex substitute the axis of symmetry value into the original equation f(0) = -(0)^2 - 2 vertex at (0, -2) intercepts f(x) = 0 then -x^2 = 2 or x^2 = -2 no x-axis intercepts as the curve is negative definite b^-4ac = -8 therefore no real roots  y -intercept x = 0 so (0.-2) domain: all real x range: all real y